# Fundamental Oscillations of Massive Boson Stars - II

### v1

@@ Appendix D Scaling for Moment of Inertia
$\omega_{1}^{ext}=\Omega-\frac{2S}{R^{3}}=\Omega\left(1-\frac{2I}{R^{3}}\right)~{}.$ (D.2)
$\omega_{1}$ is the metric function $(t,\phi)$component of the metric in the Boyer-Lindquist-type coordinates [100]. We define the scaling $I=I^{\prime}x^{3}M_{pl}^{3}$. Thus, $\Omega$ and $\omega_{1}$ remain invariant under this transformation. The complete expression for the moment of inertia used for computation is given by
@@
Using $I=I^{\prime}x^{3}M_{pl}^{3}$, the equations become independent of $x$. The dimensionless moment of inertia ($\bar{I}$) is defined as
$\bar{I}\equiv\frac{I}{(GM)^{3}}=\frac{I^{\prime}}{M^{\prime 3}}=\bar{I^{\prime}}~{}.$ (D.5)
@@
The $I$-Love-$Q$ universal relations connecting the moment of inertia, the tidal deformability, and the spin-induced quadrupole moments for NSs have been known for a long time [100]. This means that they are followed regardless of the underlying EoS. These relations were shown to be followed by dark BSs by Maselli et al. [41] for select values of $\lambda\in{0.5,1.0,1.5}\pi$ and $m\in{300,400}$ MeV. These are shown in the orange patch in Fig. 1. They found that the dark BSs with these model parameters admit the $I$-Love-$Q$ universal relations, i.e., all BS EoSs followed it regardless of the model parameters. This was confirmed by another recent work for a few BS models [87]. This was an empirical observation. Here, we prove theoretically that BSs follow these universal relations.
We focus on the $I$-Love relations here and their dependence on compactness. We do not consider the spin-induced quadrupole moments as we restrict them to non-spinning BSs. Also, spinning scalar BSs are known to lose sphericity from torus/donut-shaped configurations [35]. Note that in this relation, the dimensionless moment of inertia ($\bar{I}=I/M^{3}$) is related to the dimensionless tidal deformability ($\Lambda$). In Appendix A and Appendix. D we explicitly show that $\Lambda$ and $\bar{I}$ are independent of model parameters ($\lambda$, $m$). Thus, we can obtain the $\bar{I}-\Lambda$ curve for any EoS for BS, and this is followed by all the EoSs. This is also the reason why all the curves $I$-Love-$Q$ obtained for different EoS parameters in [41] exactly overlap without having any spread.
We show this $\bar{I}-\Lambda$ in Fig.15(a). The black dashed curve is the numerical solution obtained for the scaled EoS. We further plot the relation obtained for NSs by [100] for the sake of comparison in green. We conclude there is a degeneracy with NSs, and this relation does not break it. The blue curve shows the $I$-Love fit as obtained in [41] for BSs. The fit coefficients are given in Table. 4 with $X=\log{\Lambda}$, and $Y=\log{\bar{I}}$. This is in excellent agreement with our results as it fits within an accuracy of $1\%$, thus providing a check of veracity for our results. However, the same fit had an error of about $10\%$ in [41] (see Fig. 7 of [41]). Thus, we find the fit to be better than what was claimed before and do not provide any new relation here. [41] also find a universal relation for $k_{2}-C$. Note that [87] also reported a similar fit relation using select BS models. In Appendix. A, we have proved the same, i.e., we have shown that $k_{2}$ is also EoS independent, i.e., they do not scale with $x$.
We also show the $C$-Love and $\bar{I}$-C relations in Figs. 15(b) and 15(c), respectively. We did not find these relations in the literature for BSs. Both of these are decreasing functions. We use the same functional forms for BSs as used in [101] for NSs and quark stars (QSs). For the $C$-Love relation we use
$C=\sum_{k=0}^{4}b_{k}(\ln{\Lambda})^{k}~{},$ (E.1)
while for the $\bar{I}$-$C$ relation we use
$\bar{I}=\sum_{k=1}^{4}b_{k}C^{-k}~{}.$ (E.2)
These are all in the form Eq. 4.1. The fit coefficients for all the cases are given in Table. 4. The corresponding relations for NSs reported in [101] are shown in blue dashed curves. We plot it in only those regimes for which they were derived, i.e., for $C\geq 0.1$ and $\Lambda\leq 10^{4}$. We find that both the relations diverge with higher compactness (or higher mass, or lower $\Lambda$).
In the case of NSs, the relations involving compactness have an uncertainty of up to $10\%$ owing to the uncertainty in EoS [101], while the $I$-Love relation is much more accurate with an error under $1\%$ [101]. In the case of BS, however, we claim that all these relations are exact, i.e., the relation does not depend at all on the model parameters (or equivalently on $x$). All the fits have an error below $2\%$.
$Y$ $X$ $b_{0}$ $b_{1}$ $b_{2}$ $b_{3}$ $b_{4}$
$\log{\bar{I}}$ $\log{\Lambda}$ 1.38 9.46 ×$\times 10^{-2}$ 1.84×$\times 10^{-2}$5.84×$\times 10^{-4}$ 5.51×$\times 10^{-6}$
$C$ $\log{\Lambda}$ 2.859 ×$\times 10^{-1}$2.068 ×$\times 10^{-2}$8.276×$\times 10^{-4}$ 1.139×$\times 10^{-4}$2.781×$\times 10^{-6}$
$\bar{I}$ $1/C$- 1.532 ×$\times 10^{-1}$ 2.664 ×$\times 10^{-1}$1.317×$\times 10^{-4}$ 1.535 ×$\times 10^{-6}$
@@
Table 4: Fitting coefficients for the fits between the quantities $\bar{I^{\prime}}$, $\Lambda^{\prime}$, and $C^{\prime}$ for massive BSs in strong-interaction limit. The fitting function is given in Eqs. 4.1. The coefficients for the $I$-Love relation are the same as they were reported in [41].
@@
Figure 15: The (a) $\bar{I}$-Love (b) $C$-Love and (c) $\bar{I}$-$C$ relations for massive BSs in the strong-interaction limit. The blue curve in (a) is the fit provided in [41]. For the other two relations, we provide the fits we performed as discussed in the text. The $I$-Love-$C$ fits as provided in [101] for NSs are shown in green. We have shown this up to $C=0.1$ ($\Lambda=10^{4}$) as appropriate for NSs. The errors corresponding to the fits are shown in the lower panels in each case.

### v3

@@ Appendix B Scaling for Moment of Inertia
$\omega_{1}^{ext}=\Omega-\frac{2S}{R^{3}}=\Omega\left(1-\frac{2I}{R^{3}}\right)~.$ (B.2)
$\omega_{1}$ is the metric function $(t,\phi)$component of the metric in the Boyer-Lindquist-type coordinates [YagiYunes2013]. We define the scaling $I=I^{\prime}x^{3}M_{\rm Pl}^{3}$. Thus, $\Omega$ and $\omega_{1}$ remain invariant under this transformation. The complete expression for the moment of inertia used for computation is given by
$I$ $=\frac{8\pi}{3}\frac{1}{\Omega}\int_{0}^{R}\frac{e^{-(\hat{\nu}+\hat{\lambda})/2}R^{5}G(\rho+p)\omega_{1}}{R-2GM(R)}dR~,$ (B.3)
$\implies I$ $=\frac{8\pi}{3}\frac{1}{\Omega^{\prime}}\int_{0}^{R^{\prime}}\frac{e^{-(\hat{\nu}^{\prime}+\hat{\lambda}^{\prime})/2}R^{\prime 5}(\rho^{\prime}+p^{\prime})\omega^{\prime}_{1}(x^{3}M_{\rm Pl}^{3})}{R^{\prime}-2M^{\prime}(R^{\prime})}dR^{\prime}$ (B.4)
Here, $\hat{\nu}$ and $\hat{\lambda}$ are metric function. Using $I=I^{\prime}x^{3}M_{\rm Pl}^{3}$, the equations become independent of $x$. The dimensionless moment of inertia ($\bar{I}$) is defined as
$\bar{I}\equiv\frac{I}{(GM)^{3}}=\frac{I^{\prime}}{M^{\prime 3}}=\bar{I^{\prime}}~.$ (B.5)
Appendix C The II-LoveC-C Relations
The II-Love-$Q$ universal relations connecting the moment of inertia, the tidal deformability, and the spin-induced quadrupole moments for NSs have been known for a long time [YagiYunes2013]. This means that they are followed regardless of the underlying EoS. These relations were shown to be followed by dark BSs by Maselli et al. [Maselli2017] for select values of $\lambda\in{0.5,1.0,1.5}\pi$ and $m\in{300,400}$ MeV. These are shown in the orange patch in Fig. 1. They found that dark BSs with these model parameters admit the II-Love-$Q$ universal relations, i.e., all BS EoSs followed it regardless of the model parameters. This was confirmed by another recent work for a few BS models [Wu2023]. This was an empirical observation. Here, we prove in theory that BSs follow these universal relations. We do not consider the spin-induced quadrupole moments as we restrict them to non-spinning BSs. Note that spinning scalar BSs without interactions are known to lose sphericity from torus/donut-shaped configurations [Sanchis-Gual2019].
In Appendix A and Appendix. B we explicitly show that $\bar{\Lambda}$ and $\bar{I}$ are independent of model parameters ($\lambda$, $m$). Thus, we can obtain the $\bar{I}-\bar{\Lambda}$ curve for any BS EoS, and this will be followed by all the EoSs. This is also the reason why all the curves II-Love-$Q$ obtained for different EoS parameters in [Maselli2017] exactly overlap without having any spread. We show this $\bar{I}-\bar{\Lambda}$ in Fig.10(a). The black dashed curve is the numerical solution obtained for the scaled EoS. We further plot the fit relation obtained for NSs by [YagiYunes2013] for the sake of comparison in green. We conclude there is a degeneracy with NSs, and this relation does not break it. The blue curve shows the $I$-Love fit as obtained in [Maselli2017] for BSs (coefficients are given in Table. 4) with $X=\log{\bar{\Lambda}}$, and $Y=\log{\bar{I}}$. This fit (blue) is in excellent agreement with our numerical results as it fits within an accuracy of $1\%$, thus providing a veracity check for our results. However, the same fit had an error of about $10\%$ in [Maselli2017] (see Fig. 7 of [Maselli2017]). Thus, we find the fit to be better than what was claimed before and do not provide any new relation here. Note that [Wu2023] also reported a similar fit relation using select BS models. [Maselli2017] also found a universal relation for $k_{2}-C$. In Appendix. A, we have proved the same, i.e., we have shown that $k_{2}$ is also EoS independent, i.e., they do not scale with $x$.
We also show the $C$-Love and $\bar{I}$-C relations in Figs. 10(b) and 10(c), respectively. We did not find these relations in the literature for BSs. We use the same functional forms for BSs as used in [Yagi2017] for NSs and strange quark stars. For the $C$-Love relation we use
$C=\sum_{k=0}^{4}b_{k}(\ln{\bar{\Lambda}})^{k}~,$ (C.1)
while for the $\bar{I}$-$C$ relation we use
$\bar{I}=\sum_{k=1}^{4}b_{k}C^{-k}~.$ (C.2)
The fit coefficients for all the cases are given in Table. 4. These analytical fits to the numerically obtained relations have an error below $2\%$. The corresponding fit relations for NSs reported in [Yagi2017] are shown in green dashed curves. We plot it in only those regimes for which they were derived, i.e., for $C\geq 0.1$ and $\bar{\Lambda}\leq 10^{4}$. We find that both relations differ from the BS relation at higher compactness (or higher mass, or lower $\bar{\Lambda}$).
In the case of NSs, the relations involving compactness have an uncertainty of up to $10\%$ owing to uncertainty in the EoS [Yagi2017], while the $I$-Love relation is much more accurate for NSs with an error under $1\%$ [Yagi2017]. In the case of BSs, however, we have shown mathematically that each of these relations is a unique curve in the strong interaction limit and does not depend on the BS model parameters. Hence, these relations are exactly the same for all massive BSs in the strong interaction limit.
$Y$ $X$ $b_{0}$ $b_{1}$ $b_{2}$ $b_{3}$ $b_{4}$
@@ Appendix C The II-LoveC-C Relations
Table 4: Fitting coefficients for the fits between the quantities $\bar{I}$, $\bar{\Lambda}$, and $C$ for massive BSs in strong-interaction limit. The fitting function is given in Eqs. 4.1. The coefficients for the $\bar{I}$-Love relation are the same as they were reported in [Maselli2017].
@@
Figure 10: The (a) $\bar{I}$-Love (b) $C$-Love and (c) $\bar{I}$-$C$ relations for massive BSs in the strong-interaction limit. The blue curve in (a) is the fit provided in [Maselli2017]. For the other two relations, we provide the fits we performed as discussed in the text. The II-Love-$C$ fits as provided in [Yagi2017] for NSs are shown in green. We have shown this up to $C=0.1$ ($\bar{\Lambda}=10^{4}$) as appropriate for NSs. The errors corresponding to the fits are shown in the lower panels in each case.
Acknowledgments
L.S. and J.S.B. acknowledge support by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) through the CRC-TR 211 “Strong-interaction matter under extreme conditions” – project no. 315477589 – TRR 211. S.S. acknowledges discussions with Surhud More, with Paolo Pani at the International Centre for Theoretical Sciences (ICTS) program - Beyond the Horizon: Testing the black hole paradigm (code: ICTS/BTH2025/03), and the online workshop “Dynamical Tracers of the Nature of Dark Matter" which contributed to the improvement of the manuscript. The authors thank the anonymous referee for their comments, which helped improve the manuscript significantly and enhance the interpretation of the results.